### v1

# Persistent Homology via Finite Topological Spaces

Abstract
We propose a functorial framework for persistent homology based on finite topological spaces and their associated posets. Starting from a finite metric space, we associate a filtration of finite topologies whose structure maps are continuous identity maps. By passing functorially to posets and to simplicial complexes via crosscut constructions, we obtain persistence modules without requiring inclusion relations between the resulting complexes. We show that standard poset-level simplifications preserve persistent invariants and prove stability of the resulting persistence diagrams under perturbations of the input metric in a density-based instantiation.
2020 Mathematics Subject Classification. Primary: 55N31; Secondary: 06A11, 55U10.
Keywords: Persistent homology; finite topological spaces; posets; crosscut complexes; stability.
Bahçeşehir University, Faculty of Engineering
and Natural Sciences, Istanbul, Turkey
e-mail: selcuk.kayacan@bau.edu.tr
1 Introduction
Persistent homology has become a central tool in topological data analysis, providing multiscale descriptors of data through filtrations of simplicial complexes. Most standard constructions, such as Vietoris–Rips and Čech complexes, are built directly from metric data and rely on inclusion-based filtrations. While highly effective, these approaches often conceal order-theoretic and topological structures that arise naturally from the data and may be exploited prior to simplicialization.
In this paper, we develop a functorial framework for persistent homology that interpolates between metric data and simplicial complexes by way of finite topological spaces and posets. Starting from a finite metric space, we first associate to it a family of finite topologies indexed by a scale parameter. These topologies encode relational or density-based information and form a filtration via continuous identity maps. By passing functorially to associated posets and then to simplicial complexes via crosscut constructions, we obtain persistence modules without requiring inclusion relations between the resulting complexes.
This perspective separates the role of the metric as a generator of topological structure from the subsequent homological analysis. As a result, it allows for topological and combinatorial simplifications at the poset level, while preserving persistent invariants. Moreover, stability of the resulting persistence diagrams follows from the functoriality of the construction and elementary interleaving arguments.
Contributions
The main contributions of this paper are as follows:
  • We introduce a functorial pipeline from finite metric spaces to persistence modules via finite topological spaces and posets.
  • We show that the construction yields well-defined persistence modules even in the absence of inclusion relations between the associated simplicial complexes.
  • We employ crosscut complexes to pass from posets to simplicial complexes while preserving homotopy type.
  • We demonstrate that poset-level simplifications, such as beat-point removal, do not affect the resulting persistence diagrams.
  • We prove stability of the resulting persistence diagrams under perturbations of the input metric in a density-based instantiation.
2 Background
In this section, we shall review the basic definitions and results that are needed in this paper.
2.1 Persistent Homology
Persistent homology provides an algebraic framework for studying the evolution of topological features across a family of spaces indexed by a parameter. We briefly recall the definitions and results needed later; standard references include [4], [3], and [5].
[…]
Definition.
A persistence module indexed by $I$ is a covariant functor
[…]
Explicitly, this consists of a family of vector spaces $\{V_{i}\}_{i\in I}$ together with linear maps
$v_{i,j}:V_{i}\to V_{j}\quad\text{for all }i\leq j,$
such that $v_{i,i}=\mathrm{id}$ and $v_{j,k}\circ v_{i,j}=v_{i,k}$ whenever $i\leq j\leq k$.
Persistent homology is the study of persistence modules arising from filtrations via homology. Let $\mathbf{Top}$ denote the category of topological spaces and continuous maps.
Definition.
A filtration indexed by $I$ is a functor
$\mathbb{X}:I\longrightarrow\mathbf{Top}.$
[…]
Under suitable finiteness assumptions (e.g. pointwise finite-dimensionality), a persistence module admits a decomposition into interval modules, which can be encoded combinatorially.
Definition.
The persistence diagram $\mathsf{Dgm}(\mathbb{V})$ of a persistence module $\mathbb{V}$ is the multiset of points $(b,d)\in\overline{\mathbb{R}}^{2}$ corresponding to the birth and death parameters of its interval summands. Points on the diagonal $b=d$ are typically ignored, as they represent trivial features.
Two metrics play a central role in comparing persistence modules.
Definition.
The bottleneck distance $d_{B}$ between persistence diagrams is defined as the infimum over all matchings between diagrams of the maximum $L^{\infty}$-distance between matched points.
[…]
$\mathbb{W}(\varepsilon)(t):=\mathbb{W}(t+\varepsilon),$
with structure maps
$\mathbb{W}(\varepsilon)(t\leq s):=\mathbb{W}(t+\varepsilon\leq s+\varepsilon).$
Equivalently, $\mathbb{W}(\varepsilon)$ is obtained by precomposing $\mathbb{W}$ with the translation functor
$T_{\varepsilon}:(\mathbb{R},\leq)\longrightarrow(\mathbb{R},\leq),\qquad t\longmapsto t+\varepsilon.$
Intuitively, the shift $\mathbb{W}(\varepsilon)$ delays the appearance of all homological features by $\varepsilon$.
Definition.
Let $\varepsilon\geq 0$. Two persistence modules $\mathbb{V},\mathbb{W}$ indexed by $I\subset\mathbb{R}$ are said to be $\varepsilon$-interleaved if there exist natural transformations
$\mathbb{V}\Rightarrow\mathbb{W}(\varepsilon),\qquad\mathbb{W}\Rightarrow\mathbb{V}(\varepsilon),$
satisfying the standard commutativity conditions ensuring that compositions agree with the structure maps shifted by $2\varepsilon$. The interleaving distance $d_{I}(\mathbb{V},\mathbb{W})$ is the infimum of such $\varepsilon$.
A fundamental result in persistence theory states that the bottleneck distance $d_{B}$ and the interleaving distance $d_{I}$ coincide.
Theorem 1 (Isometry Theorem [3]).
For pointwise finite-dimensional persistence modules indexed by $\mathbb{R}$,
$d_{I}(\mathbb{V},\mathbb{W})=d_{B}(\mathsf{Dgm}(\mathbb{V}),\mathsf{Dgm}(\mathbb{W})).$
This theorem underpins the stability of persistent homology: small perturbations at the level of persistence modules lead to small changes in persistence diagrams. Finally, we recall a standard invariance principle.
Theorem 2 (Persistence Equivalence [4]).
[…]
This result allows one to replace a persistence module by any naturally isomorphic one without affecting its persistent invariants.
2.2 Finite Topological Spaces and Posets
Finite topological spaces admit a convenient combinatorial description in terms of preorders and posets. We briefly recall the basic constructions and results that will be used later; a comprehensive reference is [1].
Let $X$ be a finite topological space. For each $x\in X$, define
[…]
The collection $\{U_{x}\mid x\in X\}$ forms a minimal basis for the topology on $X$. As one can easily observe the relation $\leq$ on $X$ defined by
$x\leq y\quad\Longleftrightarrow\quad x\in U_{y}$
is a preorder. Conversely, given a finite preordered set $(X,\leq)$, one may recover a topology by declaring the sets
$U_{x}:=\{\,y\in X\mid y\leq x\,\},\qquad x\in X,$
to be a minimal basis. This establishes an equivalence between finite topological spaces and finite preorders.
Given a preorder $(X,\leq)$, define an equivalence relation by
[…]
The quotient $X/{\sim}$ inherits a natural topology whose associated preorder is a partial order.
Theorem 3 ([1, Proposition 1.3.1]).
Let $X$ be a finite topological space. Then $X$ is homotopy equivalent to the quotient space $X/{\sim}$.
In particular, every finite topological space is homotopy equivalent to a finite $T_{0}$-space, and finite $T_{0}$-spaces are in bijection with finite posets.
Let $\mathcal{P}$ be a finite poset.
Definition.
The order complex $\Delta(\mathcal{P})$ is the simplicial complex whose simplices are the finite chains in $\mathcal{P}$.
Order complexes provide a bridge between combinatorial and simplicial topology.
Theorem 4 ([1, Theorem 1.4.6]).
If $X$ is a finite $T_{0}$-space with associated poset $\mathcal{P}$, then $X$ is weakly homotopy equivalent to the order complex $\Delta(\mathcal{P})$.
Since $\Delta(\mathcal{P})$ captures the homotopy type of $X$ up to weak equivalence, the homology groups of $\Delta(\mathcal{P})$ and $X$ coincide.
While order complexes may be large, they often admit smaller simplicial models with the same homotopy type. Let $\mathcal{P}$ be a finite poset. A subset $\mathcal{C}\subseteq\mathcal{P}$ is called a crosscut if:
  1. (i)
    $\mathcal{C}$ is an antichain;
  2. (ii)
    every chain in $\mathcal{P}$ is comparable with at least one element of $\mathcal{C}$;
  3. (iii)
    for any $\mathcal{A}\subseteq\mathcal{C}$, if $\mathcal{A}$ has an upper bound or a lower bound in $\mathcal{P}$, then the join $\bigvee\mathcal{A}$ or the meet $\bigwedge\mathcal{A}$ exists in $\mathcal{P}$ (see [2] for details).
Definition.
Let $\mathcal{C}$ be a crosscut in a finite poset $\mathcal{P}$. The crosscut complex $\Delta(\mathcal{P},\mathcal{C})$ is the simplicial complex whose simplices are the subsets of $\mathcal{C}$ that admit a common upper bound or a common lower bound in $\mathcal{P}$.
Theorem 5 (Crosscut Theorem [6, Theorem 3], [2, Theorem 10.8]).
The crosscut complex $\Delta(\mathcal{P},\mathcal{C})$ is homotopy equivalent to the order complex $\Delta(\mathcal{P})$.
3 Construction
Let $(X,d)$ be a finite metric space. The metric on $X$ determines its pairwise distance matrix, which in turn encodes all geometric information about $X$ up to rigid motions and symmetries whenever $X$ arises as a finite subset of Euclidean space. This suggests that the distance matrix itself may be regarded as a primary data object, and that different constructions may extract topological information from this data at different scales or resolutions.
Our approach proceeds by associating to $(X,d)$ a family of topological spaces in a functorial manner, and then passing to combinatorial models suitable for homological analysis.
3.1 Topological Models from Metric Data
Any rule that assigns to $X$ a topology in a manner compatible with maps between metric spaces may be regarded as a topological model of the data. Given a topology $\mathcal{T}$ on $X$, we obtain a finite topological space $(X,\mathcal{T})$.
Associated to $(X,\mathcal{T})$ is its preorder defined by $x\leq y$ if and only if $x\in U_{y}$, where $U_{y}$ is the minimal basis element for the topology $\mathcal{T}$ containing $y$. Hence, one obtains a quotient space $X/{\sim}$ which is a finite $T_{0}$-space. The induced preorder on $X/{\sim}$ is in fact a partial order, yielding a finite poset $\mathcal{P}$.
From $\mathcal{P}$ we may form its order complex $\Delta(\mathcal{P})$, as well as crosscut complexes associated to suitable crosscuts.
Proposition 6.
Let $(X,\mathcal{T})$ be a finite topological space, $\mathcal{P}$ the associated poset, and $\mathcal{C}$ a crosscut consisting of maximal elements of $\mathcal{P}$. Then $(X,\mathcal{T})$ and the crosscut complex $\Delta(\mathcal{P},\mathcal{C})$ are weakly homotopy equivalent.
Proof.
By Theorem 3, $(X,\mathcal{T})$ is homotopy equivalent to its $T_{0}$-quotient. By Theorem 4, this quotient is weakly homotopy equivalent to the order complex $\Delta(\mathcal{P})$. The result now follows from the Crosscut Theorem (Theorem 5). ∎
Thus, any topology on $X$ determines a simplicial complex that is typically much smaller and more amenable to homological computation.
3.2 Filtrations and Functoriality
Let $I$ be a totally ordered finite index set. Suppose we are given a family of topologies
$\mathcal{T}_{1}\supseteq\mathcal{T}_{2}\supseteq\cdots\supseteq\mathcal{T}_{m}$
on the underlying set $X$. This defines a filtration
$\mathbb{X}\colon I\longrightarrow\mathbf{Top},$
by assigning $\mathbb{X}(i):=(X,\mathcal{T}_{i})$, and mapping each morphism $i\leq j$ in $I$ to the identity map $\mathrm{id}_{X}\colon(X,\mathcal{T}_{i})\longrightarrow(X,\mathcal{T}_{j})$, which is continuous since $\mathcal{T}_{j}\subseteq\mathcal{T}_{i}$.
Applying the constructions of Section 3.1 objectwise yields a corresponding filtration of simplicial complexes
$\mathbb{K}\colon I\longrightarrow\mathbf{Simp},\qquad i\longmapsto K_{i}:=\Delta(\mathcal{P}_{i},\mathcal{C}_{i}).$
Proposition 7.
For each $n\geq 0$, the persistence modules $H_{n}(\mathbb{X})$ and $H_{n}(\mathbb{K})$ are naturally isomorphic.
Proof.
Each step of the construction—passing from a finite topological space to its associated poset, and from a poset to a crosscut complex—is functorial with respect to continuous maps. Moreover, the weak homotopy equivalences involved are natural with respect to these maps. Applying homology yields naturally isomorphic persistence modules. In particular, the induced maps on homology commute with the structure maps of the filtrations. ∎
Remark.
Before forming the crosscut complexes, one may further simplify the posets $\mathcal{P}_{i}$ using standard techniques from poset combinatorics, such as removal of beat points or other homotopy-preserving reductions. Since these simplifications do not change the homotopy type, the resulting persistence diagrams remain unchanged by the Persistence Equivalence Theorem (Theorem 2).
Remark.
The filtration $\mathbb{K}$ does not consist of simplicial complexes related by inclusions, and therefore does not fit directly into the standard persistence homology pipeline based on matrix reduction. Nevertheless, more efficient algorithms than computing the homology of each $K_{i}$ independently are possible, exploiting the functorial structure of the construction. Developing such algorithms is an interesting direction for future work.
3.3 A Density-Based Filtration
We conclude this section by describing a concrete instance of the above framework. Let $t_{1}\leq t_{2}\leq\cdots\leq t_{m}$ be real parameters. For each $i$, let $\mathcal{O}_{i}$ be a collection of subsets of $X$ determined by a density criterion with threshold $t_{i}$, and let $\mathcal{T}_{i}$ be the topology generated by $\mathcal{O}_{i}$. The monotonicity of the density thresholds ensures that
$i\leq j\quad\Longrightarrow\quad\mathcal{T}_{j}\subseteq\mathcal{T}_{i},$
and hence defines a filtration as above.
In the extreme cases, $t_{1}$ may be chosen sufficiently small so that all co-singleton subsets are declared open, while $t_{m}$ may be chosen sufficiently large so that $\mathcal{T}_{m}$ is the trivial topology. Accordingly, the associated crosscut complexes interpolate between a discrete set of vertices and a single vertex.
Stability properties of this construction under perturbations of the metric data will be established in Section 4.
Remark.
Unlike Vietoris–Rips or Čech constructions, the simplicial complexes arising from our method are not designed to recover the ambient geometric shape of the data. In particular, a point cloud sampled from a circle need not give rise to a persistent one-dimensional homology class.
Instead, our construction summarizes statistical and relational information encoded in the density structure of the data. Persistent homology is used here as a multiscale descriptor of how dense regions interact and merge across thresholds, rather than as a tool for geometric shape reconstruction.
4 Stability
In this section we establish stability of the persistence modules arising from our construction under perturbations of the input metric. The proof relies on the functorial nature of the construction and on the theory of interleavings recalled in Section 2.1.
Let $(X,d)$ and $(X,d^{\prime})$ be two finite metric spaces on the same underlying set $X$. We say that $d^{\prime}$ is an $\varepsilon$-perturbation of $d$ if
$|d(x,y)-d^{\prime}(x,y)|\leq\varepsilon\quad\text{for all }x,y\in X.$
Such perturbations arise naturally, for example, from noise in measurements or discretization effects.
4.1 Stability of the Induced Filtrations
Fix a family of threshold values $\{t\}_{t\in\mathbb{R}}$ and let
$\mathbb{X}^{d},\mathbb{X}^{d^{\prime}}\colon(\mathbb{R},\leq)\longrightarrow\mathbf{Top}$
denote the filtrations obtained from $(X,d)$ and $(X,d^{\prime})$, respectively, via the density-based procedure described in Section 3.3.
Lemma 8.
If $d^{\prime}$ is an $\varepsilon$-perturbation of $d$, then for every $t\in\mathbb{R}$ there exist continuous identity maps
$(X,\mathcal{T}^{d}_{t})\longrightarrow(X,\mathcal{T}^{d^{\prime}}_{t+\varepsilon}),\qquad(X,\mathcal{T}^{d^{\prime}}_{t})\longrightarrow(X,\mathcal{T}^{d}_{t+\varepsilon}).$
Proof.
By construction, the topology $\mathcal{T}^{d}_{t}$ is generated by a family $\mathcal{O}^{d}_{t}$ of subsets of $X$, where membership of a subset $S\subseteq X$ in $\mathcal{O}^{d}_{t}$ depends only on whether its density (computed using $d$) exceeds the threshold $t$.
If $d^{\prime}$ is an $\varepsilon$-perturbation of $d$, then any density inequality defining membership in $\mathcal{O}^{d}_{t}$ with respect to $d$ remains valid at level $t+\varepsilon$ with respect to $d^{\prime}$. Consequently,
$\mathcal{O}^{d}_{t}\subseteq\mathcal{O}^{d^{\prime}}_{t+\varepsilon},\quad\text{and hence}\quad\mathcal{T}^{d^{\prime}}_{t+\varepsilon}\subseteq\mathcal{T}^{d}_{t}.$
This implies that the identity map on $X$ is continuous from $(X,\mathcal{T}^{d}_{t})$ to $(X,\mathcal{T}^{d^{\prime}}_{t+\varepsilon})$. The second map is obtained symmetrically. ∎
4.2 Interleaving of Persistence Modules
Applying the functorial constructions of Section 3 yields filtrations of simplicial complexes
$\mathbb{K}^{d},\mathbb{K}^{d^{\prime}}\colon(\mathbb{R},\leq)\longrightarrow\mathbf{Simp}.$
Proposition 9.
For each $n\geq 0$, the persistence modules $H_{n}(\mathbb{K}^{d})$ and $H_{n}(\mathbb{K}^{d^{\prime}})$ are $\varepsilon$-interleaved. No inclusion relations between the complexes are required.
Proof.
The identity maps of the previous lemma induce, functorially, simplicial maps
$K^{d}_{t}\longrightarrow K^{d^{\prime}}_{t+\varepsilon},\qquad K^{d^{\prime}}_{t}\longrightarrow K^{d}_{t+\varepsilon}.$
Functoriality ensures that these maps are compatible with the filtrations and satisfy the coherence conditions required for an $\varepsilon$-interleaving. Applying homology completes the proof. ∎
4.3 Stability of Persistence Diagrams
Combining the above with the Isometry Theorem yields the desired stability result.
Theorem 10 (Stability).
For each $n\geq 0$,
$d_{B}\bigl(\mathsf{Dgm}(H_{n}(\mathbb{K}^{d})),\mathsf{Dgm}(H_{n}(\mathbb{K}^{d^{\prime}}))\bigr)\leq\varepsilon.$
Proof.
By the previous proposition, the persistence modules are $\varepsilon$-interleaved. The result follows from the Isometry Theorem (Theorem 1). ∎

### v2

@@
 Starting from a finite metric space, we associate a filtration of finite topologies whose structure maps are continuous identity maps.
-By passing functorially to posets and to simplicial complexes via crosscut constructions, we obtain persistence modules without requiring inclusion relations between the resulting complexes.
-We show that standard poset-level simplifications preserve persistent invariants and prove stability of the resulting persistence diagrams under perturbations of the input metric in a density-based instantiation.
+By passing functorially to posets and to order complexes, we obtain persistence modules without requiring inclusion relations between the resulting complexes.
+We show that standard poset-level simplifications preserve persistent invariants and establish stability of the resulting persistence diagrams under perturbations of the input metric in a basic density-based instantiation, illustrating how stability arguments arise naturally in our framework.
+We further introduce a concrete density-guided construction, designed to be faithful to anchor neighborhood structure at each scale, and demonstrate its practical viability through an implementation tested on real datasets.
 2020 Mathematics Subject Classification. Primary: 55N31; Secondary: 06A11, 55U10.
-Keywords: Persistent homology; finite topological spaces; posets; crosscut complexes; stability.
+Keywords: Persistent homology; finite topological spaces; posets; order complexes; stability.
 Bahçeşehir University, Faculty of Engineering
@@
 These topologies encode relational or density-based information and form a filtration via continuous identity maps.
-By passing functorially to associated posets and then to simplicial complexes via crosscut constructions, we obtain persistence modules without requiring inclusion relations between the resulting complexes.
+By passing functorially to associated posets and then to order complexes, we obtain persistence modules without requiring inclusion relations between the resulting complexes.
 This perspective separates the role of the metric as a generator of topological structure from the subsequent homological analysis.
 As a result, it allows for topological and combinatorial simplifications at the poset level, while preserving persistent invariants.
-Moreover, stability of the resulting persistence diagrams follows from the functoriality of the construction and elementary interleaving arguments.
+Moreover, for a basic density-based instantiation of the framework, stability of the resulting persistence diagrams follows from functoriality and elementary interleaving arguments.
+This toy case serves to illustrate how stability considerations arise naturally in our setting, even though it does not yet capture informative structural features of the data.
+In particular, in the basic density-based filtration the associated posets become contractible as soon as uncovered points appear, causing the topology to collapse to a cone and limiting the descriptive power of the resulting persistence.
+This observation motivates the introduction of a more refined, density-guided construction that aims to remain faithful to anchor neighborhood structure throughout the filtration.
 Contributions
@@
   • We show that the construction yields well-defined persistence modules even in the absence of inclusion relations between the associated simplicial complexes.
-  • We employ crosscut complexes to pass from posets to simplicial complexes while preserving homotopy type.
   • We demonstrate that poset-level simplifications, such as beat-point removal, do not affect the resulting persistence diagrams.
-  • We prove stability of the resulting persistence diagrams under perturbations of the input metric in a density-based instantiation.
+  • We establish stability of the resulting persistence diagrams under perturbations of the input metric in a basic density-based instantiation, illustrating the natural compatibility of the framework with interleaving-based stability arguments.
+  • We introduce a concrete density-guided construction designed to preserve anchor neighborhood structure across scales, and provide an implementation demonstrating its applicability to real datasets.
 2 Background
@@
 Definition.
-The bottleneck distance $d_{B}$ between persistence diagrams is defined as the infimum over all matchings between diagrams of the maximum $L^{\infty}$-distance between matched points.
+The bottleneck distance $d_{B}$ between persistence diagrams is defined as the infimum over all matchings between diagrams of the maximum $\ell^{\infty}$-distance between matched points.
 […]
@@
 Finite topological spaces admit a convenient combinatorial description in terms of preorders and posets.
-We briefly recall the basic constructions and results that will be used later; a comprehensive reference is [1].
+We briefly recall the basic constructions and results that will be used later; see Barmak [1] for a comprehensive modern treatment, building on earlier work of Alexandrov and McCord.
 Let $X$ be a finite topological space. For each $x\in X$, define
@@
 Since $\Delta(\mathcal{P})$ captures the homotopy type of $X$ up to weak equivalence, the homology groups of $\Delta(\mathcal{P})$ and $X$ coincide.
-While order complexes may be large, they often admit smaller simplicial models with the same homotopy type.
-Let $\mathcal{P}$ be a finite poset.
-A subset $\mathcal{C}\subseteq\mathcal{P}$ is called a crosscut if:
-  1. (i)
-    $\mathcal{C}$ is an antichain;
-  2. (ii)
-    every chain in $\mathcal{P}$ is comparable with at least one element of $\mathcal{C}$;
-  3. (iii)
-    for any $\mathcal{A}\subseteq\mathcal{C}$, if $\mathcal{A}$ has an upper bound or a lower bound in $\mathcal{P}$, then the join $\bigvee\mathcal{A}$ or the meet $\bigwedge\mathcal{A}$ exists in $\mathcal{P}$ (see [2] for details).
-Definition.
-Let $\mathcal{C}$ be a crosscut in a finite poset $\mathcal{P}$.
-The crosscut complex $\Delta(\mathcal{P},\mathcal{C})$ is the simplicial complex whose simplices are the subsets of $\mathcal{C}$ that admit a common upper bound or a common lower bound in $\mathcal{P}$.
-Theorem 5 (Crosscut Theorem [6, Theorem 3], [2, Theorem 10.8]).
-The crosscut complex $\Delta(\mathcal{P},\mathcal{C})$ is homotopy equivalent to the order complex $\Delta(\mathcal{P})$.
 3 Construction
@@
 The induced preorder on $X/{\sim}$ is in fact a partial order, yielding a finite poset $\mathcal{P}$.
-From $\mathcal{P}$ we may form its order complex $\Delta(\mathcal{P})$, as well as crosscut complexes associated to suitable crosscuts.
-Proposition 6.
-Let $(X,\mathcal{T})$ be a finite topological space, $\mathcal{P}$ the associated poset, and $\mathcal{C}$ a crosscut consisting of maximal elements of $\mathcal{P}$.
-Then $(X,\mathcal{T})$ and the crosscut complex $\Delta(\mathcal{P},\mathcal{C})$ are weakly homotopy equivalent.
-Proof.
-By Theorem 3, $(X,\mathcal{T})$ is homotopy equivalent to its $T_{0}$-quotient.
-By Theorem 4, this quotient is weakly homotopy equivalent to the order complex $\Delta(\mathcal{P})$.
-The result now follows from the Crosscut Theorem (Theorem 5).
-∎
-Thus, any topology on $X$ determines a simplicial complex that is typically much smaller and more amenable to homological computation.
+From $\mathcal{P}$ we may form its order complex $\Delta(\mathcal{P})$.
+By Theorem 3 and Theorem 4, the finite topological space $(X,\mathcal{T})$ is weakly homotopy equivalent to the order complex $\Delta(\mathcal{P})$.
+Thus, any topology on $X$ determines a simplicial complex in a canonical way and standard computational tools for simplicial complexes become available for homology calculations.
 3.2 Filtrations and Functoriality
@@
 Applying the constructions of Section 3.1 objectwise yields a corresponding filtration of simplicial complexes
-$\mathbb{K}\colon I\longrightarrow\mathbf{Simp},\qquad i\longmapsto K_{i}:=\Delta(\mathcal{P}_{i},\mathcal{C}_{i}).$
+$\mathbb{K}\colon I\longrightarrow\mathbf{Simp},\qquad i\longmapsto K_{i}:=\Delta(\mathcal{P}_{i}).$
 Proposition 5.
@@
 Proof.
-Each step of the construction—passing from a finite topological space to its associated poset, and from a poset to a crosscut complex—is functorial with respect to continuous maps.
+Each step of the construction—passing from a finite topological space to its associated poset, and from a poset to its order complex—is functorial with respect to continuous maps.
 Moreover, the weak homotopy equivalences involved are natural with respect to these maps.
@@
 Remark.
-Before forming the crosscut complexes, one may further simplify the posets $\mathcal{P}_{i}$ using standard techniques from poset combinatorics, such as removal of beat points or other homotopy-preserving reductions.
+Before forming the order complexes, one may further simplify the posets $\mathcal{P}_{i}$ using standard techniques from poset combinatorics, such as removal of beat points or other homotopy-preserving reductions.
 Since these simplifications do not change the homotopy type, the resulting persistence diagrams remain unchanged by the Persistence Equivalence Theorem (Theorem 2).
@@
 Developing such algorithms is an interesting direction for future work.
-3.3 A Density-Based Filtration
-We conclude this section by describing a concrete instance of the above framework.
+3.3 Density-Based Filtration – Basic Model
+We conclude this section by describing a toy case application of the above framework.
 Let $t_{1}\leq t_{2}\leq\cdots\leq t_{m}$ be real parameters.
@@
 In the extreme cases, $t_{1}$ may be chosen sufficiently small so that all co-singleton subsets are declared open, while $t_{m}$ may be chosen sufficiently large so that $\mathcal{T}_{m}$ is the trivial topology.
-Accordingly, the associated crosscut complexes interpolate between a discrete set of vertices and a single vertex.
-Stability properties of this construction under perturbations of the metric data will be established in Section 4.
+Accordingly, the associated order complexes interpolate between a discrete set of vertices and a single vertex.
+Despite its conceptual clarity, this basic construction has a significant limitation: as soon as uncovered points (i.e.
+points not contained in any of the chosen generating neighborhoods at that scale) appear, the associated posets acquire a maximal element, causing the topology to become contractible.
+Consequently, the filtration rapidly loses higher-order information about the data.
+While this behavior makes the basic model unsuitable for detailed data analysis, it serves as a useful conceptual setting in which functoriality and stability can be examined transparently.
+Stability properties of this basic construction under perturbations of the metric data will be established in Section 4, serving to exemplify how stability arguments arise naturally within our functorial framework.
 Remark.
@@
 4 Stability
-In this section we establish stability of the persistence modules arising from our construction under perturbations of the input metric.
+In this section we establish stability of the persistence modules arising from the basic density-based construction under perturbations of the input metric.
 The proof relies on the functorial nature of the construction and on the theory of interleavings recalled in Section 2.1.
@@
 By the previous proposition, the persistence modules are $\varepsilon$-interleaved. The result follows from the Isometry Theorem (Theorem 1). ∎
+5 Density-Guided Instance
+The basic density-based filtration introduced in Section 3.3 serves as a conceptual illustration of the functorial framework and its stability properties.
+As discussed there, however, that construction rapidly becomes topologically trivial once the topology admits a maximal element, limiting its descriptive power for real data.
+In this section, we introduce a concrete density-guided construction designed to overcome this limitation and report on a preliminary empirical study based on an explicit implementation.
+The purpose of this section is not to provide theoretical guarantees or to interpret empirical results in depth, but rather to demonstrate that the proposed framework admits practical instantiations that produce informative persistence modules on real datasets.
+5.1 The density-guided construction
+Let $(X,d)$ be a finite metric space. As before, we consider a finite sequence of density thresholds
+$\tau_{1}\leq\tau_{2}\leq\cdots\leq\tau_{m},$
+and associate to each $\tau_{i}$ a finite topology $\mathcal{T}_{i}$ on the underlying set $X$.
+In the basic density-based construction of Section 3.3, the topology $\mathcal{T}_{i}$ is generated by declaring all subsets of $X$ satisfying a density condition at scale $\tau_{i}$ to be open.
+While conceptually simple, this approach is computationally infeasible in practice, as the number of such dense subsets grows exponentially with $|X|$.
+The density-guided construction replaces this exhaustive approach with a structured and tractable generator set. We begin by fixing a finite collection of candidate anchor neighborhoods
+$\mathcal{A}=\{A_{1},\dots,A_{k}\},$
+determined once from the metric data. These anchor neighborhoods serve as potential generators of the topology across all scales.
+For each threshold $\tau_{i}$, only a subset
+$\mathcal{A}_{i}\subseteq\mathcal{A}$
+of anchor neighborhoods satisfies the prescribed density criterion at scale $\tau_{i}$. As the density threshold increases, fewer anchor neighborhoods remain eligible, yielding a nested sequence
+$\mathcal{A}_{1}\supseteq\mathcal{A}_{2}\supseteq\cdots\supseteq\mathcal{A}_{m}.$
+To control the evolution of the topology and preserve relational information between dense regions, we augment the eligible anchor neighborhoods $\mathcal{A}_{i}$ by a secondary collection of sets $\mathcal{B}_{i}$.
+The sets in $\mathcal{B}_{i}$ are not density generators themselves, but are introduced to ensure compatibility of the topology across successive scales and to prevent loss of structural information when anchor neighborhoods become ineligible.
+Their construction is constrained by the topology at the previous scale and by the requirement that the resulting topology remain faithful to the anchor structure.
+The topology $\mathcal{T}_{i}$ is then generated by the family
+$\mathcal{G}_{i}:=\mathcal{A}_{i}\cup\mathcal{B}_{i}.$
+Remark.
+While the explicit construction of the sets $\mathcal{B}_{i}$ involves several layers and is deferred to the implementation, their role is purely structural: they serve to mediate transitions between successive anchor configurations without introducing new density-driven features.
+The defining feature of the construction is that $\mathcal{T}_{i}$ is faithful to anchor structure in the following sense: every minimal nonempty intersection of anchor neighborhoods in $\mathcal{A}_{i}$ appears as an element of the minimal basis of $\mathcal{T}_{i}$.
+Equivalently, the collection of minimal nonempty intersections generated by $\mathcal{A}_{i}$ is contained in the minimal basis of the topology generated by $\mathcal{G}_{i}$.
+As the density threshold increases, the shrinking of the eligible anchor set $\mathcal{A}_{i}$ drives the evolution of the topology across scales and prevents premature collapse to contractible structures.
+The resulting family of topologies
+$\mathcal{T}_{1}\supseteq\mathcal{T}_{2}\supseteq\cdots\supseteq\mathcal{T}_{m}$
+forms a filtration via continuous identity maps, to which the functorial constructions of Sections 3.1 and 3.2 apply.
+Remark.
+The density-guided construction may thus be viewed as a filtration obtained by progressively invalidating anchor neighborhoods rather than introducing new generators.
+5.2 Implementation
+The density-guided construction described above has been implemented in a research prototype software pipeline.
+The implementation follows the functorial structure described in Sections 3.1 and 3.2 and produces, for each scale parameter, the associated finite topology, its corresponding $T_{0}$-poset, and the resulting order complex.
+Homology is computed on the $2$-skeleton of the order complex, which is sufficient for dimensions 0 and $1$.
+In addition to computing persistence modules, the implementation generates detailed diagnostic outputs at each filtration step, including per-scale summaries of topological and combinatorial quantities, persistence barcodes in both visual and machine-readable formats, and auxiliary plots for inspection.
+These outputs are included directly in the repository so that the empirical behavior of the construction can be examined without re-running the experiments.
+The implementation used for the empirical results reported in this section is publicly available [4].
+5.3 Datasets
+We evaluated the implementation on two datasets commonly used in topological data analysis and related areas:
+  • COIL-20, a collection of grayscale images of objects observed from multiple viewpoints [5].
+  • Paul15, a high-dimensional dataset frequently used in empirical studies of persistent homology [7].
+These datasets were chosen solely to verify that the construction produces well-defined filtrations and computable persistence modules on real data, and not for comparative evaluation.
+In addition, the repository includes several synthetic examples designed to illustrate and probe the behavior of the density-guided construction under controlled conditions.
+These examples are intended to provide intuition about the evolution of the associated topologies and posets across density scales, and to serve as diagnostic test cases for the implementation.
+5.4 Empirical observations
+For both datasets, the density-guided construction produced nontrivial filtrations of finite topological spaces whose associated posets retained meaningful structure across multiple scales.
+In particular, the resulting order complexes did not collapse immediately to contractible spaces as the density threshold increased.
+Using the implementation, we compute summary tables and persistence diagrams for homological dimensions 0 and $1$, documenting the evolution of connected components and one-dimensional features across density scales.
+The repository includes per-step filtration summaries (recording, among other quantities, Betti numbers, numbers of anchors, coverage statistics, and combinatorial sizes of the associated complexes), as well as barcode visualizations and machine-readable barcode data.
+These results are presented descriptively and without further interpretation.
+No claims are made regarding semantic interpretation of features or comparative performance; the purpose of this section is solely to document the empirical behavior of the construction.
+5.5 Remarks
+We conclude this section with several remarks intended to place the density-guided construction in context.
+First, while density plays a central role in many data analysis methods, persistent homology pipelines in which the scale parameter itself is driven by density appear to be relatively unexplored.
+The present framework provides a natural setting for studying how dense regions interact and merge across density levels.
+Second, the construction is not intended for geometric shape recovery.
+Instead, it aims to capture relational and statistical structure across density scales, a task that is orthogonal to classical goals in persistent homology but of independent interest.
+Finally, an advantage of the proposed approach is that homological features can be readily traced back to subsets of data points through the underlying poset structure.
+This makes it straightforward to identify the supports of homology classes, a task that is often nontrivial in standard matrix-reduction-based pipelines.
+A detailed investigation of computational aspects and theoretical properties of this feature is left for future work.